# Automatically Analyzing Performance Issues in Android Apps: How Far Are We?

### v2

@@ 2.2.1. Performance Issues in Real-world
Energy Consumption Resource Updating unnecessary sensor data in the background  (Liu2014GreenDroidAD, ; Liu2013WhereHM, ; Li2019DetectingAD, ; Behrouz2016EnergyawareTM, ; Li2022CombattingEI, ; Liu2016FixingRL, ; Couto2020EnergyRF, ; Jabbarvand2020AutomatedCO, ; Li2020EnergyDxDE, ; Guo2013CharacterizingAD, ; Lu2022DetectingRU, ; Banerjee2018EnergyPatchRR, ; Wu2016LightWeightIA, ; Jabbarvand2019SearchBasedET, ; Banerjee2014DetectingEB, ; Yan2013SystematicTF, ; Jabbarvand2017DroidAE, ; Nikzad2014APEAA, ; Wu2016Relda2AE, ; Vsquez2015HowDD, ; Bhatt2020AutomatedRO, ; Rua2023ALE, ; Liu2019DroidLeaksAC, )
Background Thread Setting background services at unnecessarily high frequencies (Liu2014CharacterizingAD, ; Banerjee2018EnergyPatchRR, ; Li2019DetectingAD, ; Banerjee2014DetectingEB, ; Jabbarvand2019SearchBasedET, )
@@
Loop Executing energy-intensive behaviors in loop  (Banerjee2014DetectingEB, ; Behrouz2016EnergyawareTM, ; Lyu2018RemoveRF, ; Fratantonio2015CLAPPCL, ; Lyu2017AnES, ; Jabbarvand2017DroidAE, ; Li2014AnES, ; Palomba2019OnTI, )
Redundant Frames Generating frames even when there are no visual changes  (Li2019DetectingAD, ; Kim2016ContentCentricEM, )
Multi-Threads Activating/deactivating a UI component simultaneously in two different threads (Pathak2012WhatIK, )
HTTP Request Dividing a large packet into small pieces and transmitting data frequently (Li2016AutomatedEO, ; Li2014AnES, )
Obsolete Task Downloading unnecessary content when switching activities  (Kang2016DiagDroidAP, ; Couto2020EnergyRF, )
Image Decoding identical image multiple times (Song2021IMGDroidDI, ; Li2019CharacterizingAD, ; Carette2017InvestigatingTE, ; Kang2016ExperienceRD, )
Service Creating or stopping services at an inappropriate timeslot  (Song2019ServDroidDS, )
Wakelock Forgetting to release wakelocks after usage  (Liu2014GreenDroidAD, ; Pathak2012WhatIK, ; Liu2016UnderstandingAD, ; Li2022CombattingEI, ; Couto2020EnergyRF, ; Li2020EnergyDxDE, ; Behrouz2016EnergyawareTM, ; Banerjee2018EnergyPatchRR, ; Banerjee2014DetectingEB, ; Pathak2012WhereIT, ; Jabbarvand2017DroidAE, ; Vsquez2015HowDD, ; Das2020CharacterizingTE, )
@@
UI Rendering 1) Views are rendered but not visible to users because they are off-screen or set to invisible  (Gao2017EveryPC, ; Lin2024AgingOG, ; Liu2014CharacterizingAD, ; Habchi2019TheRO, ; Janssen2022OnTI, )
AsyncTask Updating destroyed components during AsyncTask life-cycle (Lin2015StudyAR, ; Kang2016DiagDroidAP, ; Pan2020StaticAC, ; Lin2014RetrofittingCF, ; Fan2018EfficientlyMA, )
Inefficient Method Calling unnecessary methods (Carette2017InvestigatingTE, ; Couto2020EnergyRF, ; Habchi2019TheRO, ; Palomba2019OnTI, )
Display Screen The OLED display consumes more energy when displaying light colors than dark ones (Li2014MakingWA, ; LinaresVsquez2018MultiObjectiveOO, ; Vsquez2015OptimizingEC, ; Wan2015DetectingDE, )
@@
Advertisements Frequent advertisement and analytics (A&A) service requests lead to energy waste (Cito2016BatteryawareTI, ; Petalotis2024AnES, )
@@
View Drawing all items in a ListView without reusing existing data  (Habchi2019TheRO, ; Couto2020EnergyRF, )
CPU Usage Multi-Threads The music player mishandles metadata extraction, spawning thousands of threads, overloading the CPU.
@@
UI Rendering Spin animation has unnecessary frequent refreshes, forcing GPU rendering on every frame (Lee2019ImprovingEE, )
@@ 2.2.2. Root Causes of Performance Issues
Here, we remind readers that during the process of manual summarization, we found that the descriptions of performance issues are chaotic. The absence of a universally accepted taxonomy results in many discussions targeting the same performance problems but using different descriptive terms (Hort2021ASO, ; Liu2014CharacterizingAD, ; Habchi2019TheRO, ; Rua2023ALE, ; Das2020CharacterizingTE, ). Additionally, since one factor can lead to multiple performance issues consequences, and one consequence can be caused by multiple factors, it complicates the creation of a clear and comprehensible performance classification system. To address this gap, our work distinguishes between the consequences and factors of performance issues. Specifically, we first classify performance issues by their consequences, then define a series of factors that cause these issues, and finally link the consequences to their corresponding factors. To the best of our knowledge, we are the first to systematically classify all Android performance issues in a hierarchical manner. We now elaborate on the seven major categories, each with the corresponding 49 factors along with an example presented in Table 1.
@@
Memory Consumption. As discussed in Section 2.2.1, memory consumption is the most frequently highlighted performance issue among developers. Our real-world analysis revealed 20 factors that contribute to these memory consumption issues, which are listed under the Memory Consumption category in Table 1. These issues arise from a variety of factors, often resulting from inefficient coding practices that lead to excessive or unnecessary memory use. For example, repeatedly reading the same value into an array within an infinite loop can cause memory to be continuously occupied without being released, ultimately leading to an out-of-memory (OOM) error. Additionally, failing to destroy an AsyncTask when it holds references to active GUI components can prevent the garbage collection of the destroyed GUI, leading to memory leaks.
@@
Responsiveness. Another important aspect of Android performance issues is responsiveness. Poor responsiveness leads to perceptible delays such as image rendering delays, UI freezes, unresponsiveness, etc. In our analysis, we identified 14 key factors contributing to this issue. Many real-world scenarios show that heavy operations associated with UI threads can slow responsiveness, such as decoding large images in the UI threads. Besides, inefficient code logic can also slow down the response time. For example, improper stacking and layering in layouts can cause pixels to be redrawn multiple times within the same frame, slowing down the response time.
@@
Storage Consumption. As the third most concerning performance issue for developers, inefficient storage consumption often leads to apps occupying more storage space than expected, which can limit the device’s ability to install other apps and store data. Through our real-world exploration, we identified four key factors contributing to storage consumption issues. These factors typically arise from inefficient coding practices and poor data management, resulting in excessive or unnecessary storage usage. For instance, creating a large number of similar functional pages and activities that could have been reused, or including redundant code that never gets executed during runtime, wastes storage space. Moreover, adding unnecessary third-party library dependencies to the Gradle file further increases the app’s size after building, leading to larger APK files.
@@
Energy Consumption. As outlined in Section 2.2.1, energy consumption ranks as the second highest concern among users, as apps that drain battery faster than expected can frustrate users, especially given the limited battery resources on mobile devices. However, developers have not prioritized addressing this issue to the same extent. Consequently, in our real-world analysis, we identified only two key factors contributing to energy consumption problems. The root causes generally stem from inefficient resource usage and unnecessary background operations that provide no tangible benefits to users. For instance, updating unneeded sensor data in the background or running background services at unnecessarily high frequencies both result in excessive energy consumption. It is worth noting that, although developers do not prioritize energy efficiency, many research papers have extensively studied this issue, bridging the gap in developer attention. We will further discuss the performance issues targeted by researchers in Section 3.2.1.
@@
CPU Usage. Excessive CPU usage can cause the device to overheat, triggering the thermal management system to automatically lower the processor’s frequency to prevent overheating. This leads to a decline in overall system performance, affecting the normal operation of other apps and background services, and may even result in crashes, while also shortening the lifespan of the hardware and battery. Based on our analysis, we identified four distinct factors contributing to CPU usage issues. These issues typically stem from unnecessary or inefficient operations that waste CPU resources. For instance, using transparency in canvases where it’s not needed forces the browser to perform extra work, increasing CPU load. Similarly, auto-playing GIFs by sequentially displaying images instead of using a static preview image results in unnecessary CPU consumption.
@@
Internet Data Usage. Network resources on mobile devices are limited, especially when Wi-Fi is unavailable, requiring users to purchase cellular data. Inefficient data usage can lead to unnecessary expenses and create a negative user experience. From our analysis, we identified three key factors contributing to this issue. The root causes often involve the unnecessary consumption of network resources. For example, downloading unwanted content without user consent or automatically downloading embedded advertisements in the background can lead to excessive data usage.
Answers to RQ1.2: Performance issues are inherently complex, often arising from an interplay of multiple contributing factors, where a single issue may stem from various causes, and, conversely, a single factor may amplify multiple issues. We identified 49 distinct factors contributing to the seven categories of performance issues.
@@ 2.2.3. Common Code Patterns in Real-World
After identifying key factors of performance issues, we analyzed Stack Overflow discussions, GitHub issues, and commits to uncover common code patterns associated with these issues. We summarized six types of common patterns as follows:
@@
API Misuse. API misuse presents a significant concern in relation to performance issues, encompassing misuses such as incorrect API invocation, misuse of API orders, and misuse of API parameters. These errors can result in severe performance consequences, such as ANR and delays. For example, passing a null context (a bad practice in parameter value passing) to PendingIntent results in an Application Not Responding error. Besides, another common pattern involves the misuse of API order, as demonstrated in the code snippet 1 from an app com.mattermost.rnbeta, where dismissBottomSheet (line 4) is invoked before Clipboard.setString(postMessage) (line 5). In this scenario, Clipboard.setString has to wait until dismissBottomSheet is executed, leading to a delay in response time.

### v3

@@ 2.2.1. Performance Issues in Real-world
Energy Consumption Resource Updating unnecessary sensor data in the background  (Liu2014GreenDroidAD, ; Liu2013WhereHM, ; Behrouz2016EnergyawareTM, ; Li2022CombattingEI, ; Liu2016FixingRL, ; Couto2020EnergyRF, ; Jabbarvand2020AutomatedCO, ; Li2020EnergyDxDE, ; Guo2013CharacterizingAD, ; Lu2022DetectingRU, ; Banerjee2018EnergyPatchRR, ; Wu2016LightWeightIA, ; Jabbarvand2019SearchBasedET, ; Banerjee2014DetectingEB, ; Yan2013SystematicTF, ; Jabbarvand2017DroidAE, ; Nikzad2014APEAA, ; Wu2016Relda2AE, ; Vsquez2015HowDD, ; Bhatt2020AutomatedRO, ; Rua2023ALE, ; Liu2019DroidLeaksAC, ; Hort2021ASO, ; LeGoaer2022ecoCodeAS, ; Cruz2019CatalogOE, ; BernalCrdenas2015ImprovingEC, ; Goar2020EnforcingGC, ; Vsquez2017GEMMAMO, ; Morales2018EARMOAE, ; Das2016AQA, )
Background Thread Setting background services at unnecessarily high frequencies (Liu2014CharacterizingAD, ; Banerjee2018EnergyPatchRR, ; Li2019DetectingAD, ; Banerjee2014DetectingEB, ; Jabbarvand2019SearchBasedET, ; BernalCrdenas2015ImprovingEC, )
Wakelock Forgetting to release wakelocks after usage  (Liu2014GreenDroidAD, ; Pathak2012WhatIK, ; Liu2016UnderstandingAD, ; Li2022CombattingEI, ; Couto2020EnergyRF, ; Li2020EnergyDxDE, ; Behrouz2016EnergyawareTM, ; Banerjee2018EnergyPatchRR, ; Banerjee2014DetectingEB, ; Pathak2012WhereIT, ; Jabbarvand2017DroidAE, ; Vsquez2015HowDD, ; Das2020CharacterizingTE, ; LeGoaer2022ecoCodeAS, ; Cruz2019CatalogOE, ; Goar2020EnforcingGC, ; Iannone2020RefactoringAE, ; Vsquez2017GEMMAMO, ; Prestat2024DynAMICSAT, ; Palomba2017LightweightDO, ; MazueraRozo2020InvestigatingTA, ; Wu2023ASL, )
HTTP Request Dividing a large packet into small pieces and transmitting data frequently (Li2016AutomatedEO, ; Li2014AnES, ; Goar2020EnforcingGC, ; Vsquez2017GEMMAMO, ; Das2016AQA, ; Wu2023ASL, )
Loop Executing energy-intensive behaviors in loop  (Banerjee2014DetectingEB, ; Behrouz2016EnergyawareTM, ; Lyu2018RemoveRF, ; Fratantonio2015CLAPPCL, ; Lyu2017AnES, ; Jabbarvand2017DroidAE, ; Li2014AnES, ; Palomba2019OnTI, ; Goar2020EnforcingGC, ; Das2016AQA, )
Service Creating or stopping services at an inappropriate timeslot  (Song2019ServDroidDS, ; BernalCrdenas2015ImprovingEC, ; Goar2020EnforcingGC, ; Das2016AQA, )
Redundant Frames Generating frames even when there are no visual changes  (Li2019DetectingAD, ; Kim2016ContentCentricEM, ; LeGoaer2022ecoCodeAS, )
@@
Multi-Threads Activating/deactivating a UI component simultaneously in two different threads (Pathak2012WhatIK, ; Feng2019LearningPO, ; Das2016AQA, )
Inefficient Method Calling unnecessary methods (Carette2017InvestigatingTE, ; Couto2020EnergyRF, ; Habchi2019TheRO, ; Palomba2019OnTI, ; Vsquez2014MiningEA, ; Goar2020EnforcingGC, ; Iannone2020RefactoringAE, ; Vsquez2017GEMMAMO, ; Bangash2023CosteffectiveSF, ; Prestat2024DynAMICSAT, ; Morales2018EARMOAE, ; Palomba2017LightweightDO, ; Feng2019LearningPO, ; Das2016AQA, ; Wu2023ASL, )
Obsolete Task Downloading unnecessary content when switching activities  (Kang2016DiagDroidAP, ; Couto2020EnergyRF, ; Cruz2019CatalogOE, )
Image Decoding identical image in multiple times (Song2021IMGDroidDI, ; Li2019CharacterizingAD, ; Carette2017InvestigatingTE, ; Kang2016ExperienceRD, ; Das2016AQA, )
@@
UI Rendering Loading CSS for below-the-fold content during webpage loading wastes energy  (Gao2017EveryPC, ; Lin2024AgingOG, ; Liu2014CharacterizingAD, ; Habchi2019TheRO, ; Janssen2022OnTI, ; Cruz2019CatalogOE, ; Wu2023ASL, )
AsyncTask Updating destroyed components during AsyncTask life-cycle (Lin2015StudyAR, ; Kang2016DiagDroidAP, ; Pan2020StaticAC, ; Lin2014RetrofittingCF, ; Fan2018EfficientlyMA, ; BernalCrdenas2015ImprovingEC, ; Iannone2020RefactoringAE, )
Display Screen The OLED display consumes more energy when displaying light colors than dark ones (Li2014MakingWA, ; LinaresVsquez2018MultiObjectiveOO, ; Vsquez2015OptimizingEC, ; Wan2015DetectingDE, ; Cruz2019CatalogOE, ; BernalCrdenas2015ImprovingEC, ; Goar2020EnforcingGC, ; Vsquez2017GEMMAMO, )
@@
Advertisements Frequent advertisement and analytics (A&A) service requests lead to energy waste (Cito2016BatteryawareTI, ; Petalotis2024AnES, ; Vsquez2017GEMMAMO, )
@@
View Drawing all items in a ListView without reusing exising data  (Habchi2019TheRO, ; Couto2020EnergyRF, ; Feng2019LearningPO, )
CPU Usage Multi-Threads The music player mishandles metadata extraction, spawning thousands of threads, overloading the CPU
@@
Third-Party Library Third-party library’s bug causes a failure to properly release the resources used to display GIFs
Background Thread Frequently perform read and write operations in a background thread
Resource Use high sampling rate sensors such as the gyroscope
@@
UI Rendering Spin animation has unnecessary frequent refreshes, forcing GPU rendering on every frame (Lee2019ImprovingEE, ; Habchi2020AndroidCS, ; Wu2023ASL, )
@@
HTTP Request High-speed movement triggers frequent GPS updates, causing redundant API uploads
@@ 2.2.2. Root Causes of Performance Issues
Here, we remind readers that during the process of manual summarization, we found that the descriptions of performance issues are chaotic. The absence of a universally accepted taxonomy results in many discussions targeting the same performance problems but using different descriptive terms (Hort2021ASO, ; Liu2014CharacterizingAD, ; Habchi2019TheRO, ; Rua2023ALE, ; Das2020CharacterizingTE, ). Additionally, since one factor can lead to multiple performance issue consequences, and one consequence can be caused by multiple factors, it complicates the creation of a clear and comprehensible performance classification system. To address this gap, our work distinguishes between the consequences and factors of performance issues.
Specifically, we applied the thematic analysis methodology (Cruzes2011RecommendedSF, ; Das2016AQA, ) to analyze all the real-world data we collected, including GitHub issues and commits, Google Play Store user comments, and Stack Overflow posts. Using this approach, we classified performance issues according to their observed consequences, identified the underlying contributing factors, and established explicit links between them to construct a unified cross-source taxonomy.
This process involved the following steps:
  • (i) Familiarization with the data: Two authors independently reviewed all the real-world records to gain a thorough understanding of the content.
  • (ii) Initial coding: After familiarization, the two authors conducted open labeling by assigning descriptive tags to each record. In total, they analyzed 114 Google Play user reviews, 1,484 Stack Overflow question threads, 69 GitHub issues, and 222 GitHub commits, as described in Section 2.1.3. Each fragment was labeled independently by both authors. At this stage, we identified 226 distinct labels. For example, a review, GitHub issue, Stack Overflow post, or commit might be annotated with tags such as memory usage inefficiency, memory leak, out of memory, image decoding without resizing, or image decoding in the main thread.
  • (iii) Categorization of issues and factors: The initial labels were then organized into higher-level performance issues and contributing factors. For instance, memory usage inefficiency, memory leak, and out of memory were grouped under the broader issue category Memory Consumption. In contrast, tags such as decoding images without resizing, inefficient image format, and decoding identical images multiple times were consolidated into the broader Image factor. At this stage, we obtained 15 categories of performance issues and 96 categories of contributing factors.
  • (iv) Review and refinement: All authors collaboratively reviewed and refined the labeled data to ensure accurate classification. During this iterative process, we held multiple discussions to merge semantically similar items, split overly broad categories, and discard irrelevant or redundant tags. For example, we combined the factors Drawing View and Bottom Sheet View into the broader View category. Finally, through this systematic and iterative process, we developed a taxonomy of seven major performance issues and 63 contributing factors. An illustrative example is presented in Table 1, where the Real-World column is marked with (✓).
@@
Memory Consumption. As discussed in Section 2.2.1, memory consumption is the most frequently highlighted performance issue among developers. Our real-world analysis revealed 21 factors that contribute to these memory consumption issues, which are listed under the Memory Consumption category in Table 1. These issues arise from a variety of factors, often resulting from inefficient coding practices that lead to excessive or unnecessary memory use. For example, repeatedly reading the same value into an array within an infinite loop can cause memory to be continuously occupied without being released, ultimately leading to an out-of-memory (OOM) error. Additionally, failing to destroy an AsyncTask when it holds references to active GUI components can prevent the garbage collection of the destroyed GUI, leading to memory leaks.
@@
Responsiveness. Another important aspect of Android performance issues is responsiveness. Poor responsiveness leads to perceptible delays such as image rendering delays, UI freezes, unresponsiveness, etc. In our analysis, we identified 16 key factors contributing to this issue. Many real-world scenarios show that heavy operations associated with UI threads can slow responsiveness, such as decoding large images in the UI threads. Besides, inefficient code logic can also slow down the response time. For example, improper stacking and layering in layouts can cause pixels to be redrawn multiple times within the same frame, slowing down the response time.
@@
Storage Consumption. As the third most concerning performance issue for developers, inefficient storage consumption often leads to apps occupying more storage space than expected, which can limit the device’s ability to install other apps and store data. Through our exploration, we identified seven key factors contributing to storage consumption issues. These factors typically arise from inefficient coding practices and poor data management, resulting in excessive or unnecessary storage usage. For instance, creating a large number of similar functional pages and activities that could have been reused, or including redundant code that never gets executed during runtime, wastes storage space. Moreover, adding unnecessary third-party library dependencies to the Gradle file further increases the app’s size after building, leading to larger APK files.
@@
Energy Consumption. As outlined in Section 2.2.1, energy consumption ranks as the second highest concern among users, as apps that drain battery faster than expected can frustrate users, especially given the limited battery resources on mobile devices. However, developers have not prioritized addressing this issue to the same extent. Consequently, in our real-world analysis, we identified only six key factors contributing to energy consumption problems. The root causes generally stem from inefficient resource usage and unnecessary background operations that provide no tangible benefits to users. For instance, updating unneeded sensor data in the background or running background services at unnecessarily high frequencies both result in excessive energy consumption. It is worth noting that, although developers do not prioritize energy efficiency, many research papers have extensively studied this issue, bridging the gap in developer attention. We will further discuss the performance issues targeted by researchers in Section 3.2.1.
@@
CPU Usage. Excessive CPU usage can cause the device to overheat, triggering the thermal management system to automatically lower the processor’s frequency to prevent overheating. This leads to a decline in overall system performance, affecting the normal operation of other apps and background services, and may even result in crashes, while also shortening the lifespan of the hardware and battery. Based on our analysis, we identified seven distinct factors contributing to CPU usage issues. These issues typically stem from unnecessary or inefficient operations that waste CPU resources. For instance, using transparency in canvases where it’s not needed forces the browser to perform extra work, increasing CPU load. Similarly, auto-playing GIFs by sequentially displaying images instead of using a static preview image results in unnecessary CPU consumption.
@@
Internet Data Usage. This category primarily refers to the unnecessary or unauthorized use of network resources, regardless of whether the device is connected via Wi-Fi or cellular data, and regardless of whether the network behavior is caused by the app’s own logic or by integrated third-party APIs. Although such issues may occur under various network conditions, their negative impact is often more significant when users are not connected to Wi-Fi, as inefficient data usage can lead to unexpected expenses. From our analysis, we identified four key factors contributing to this issue. The root causes often involve unnecessary or unauthorized consumption of network resources, regardless of whether the device is connected via Wi-Fi or cellular data, and irrespective of whether the behavior is initiated by the app itself or by third-party libraries (e.g., advertising SDKs). For example, downloading unwanted content without user consent or automatically downloading embedded advertisements in the background can lead to excessive data usage.
Answers to RQ1.2: Performance issues are inherently complex, often arising from an interplay of multiple contributing factors, where a single issue may stem from various causes, and, conversely, a single factor may amplify multiple issues. We identified 63 distinct factors contributing to the seven categories of performance issues.
@@ 2.2.3. Common Code Patterns in Real-World
After identifying the key contributing factors of performance issues, we analyzed Stack Overflow discussions, GitHub issues, and commits to uncover common code patterns associated with these issues. We found that the relationships among code patterns, performance issues, and contributing factors are inherently many-to-many: a single code pattern may involve multiple contributing factors and lead to various types of performance issues; similarly, a single factor may appear across multiple different code patterns. It is worth noting that a number of prior studies have also focused on identifying and summarizing performance-related code patterns in Android applications (Das2020CharacterizingTE, ; cruz2018using, ; Liu2014CharacterizingAD, ; Nistor2014SunCatHD, ; Guo2013CharacterizingAD, ; Song2021IMGDroidDI, ; Habchi2019TheRO, ; Li2019CharacterizingAD, ; Liu2016UnderstandingAD, ; Cruz2019CatalogOE, ). For example, the well-known DrawAllocation issue typically refers to performing memory allocation within UI rendering methods such as onDraw(), and has been widely discussed in the literature (Das2020CharacterizingTE, ; cruz2018using, ) and flagged by the Android Lint tool. This specific issue often involves contributing factors such as Loop, Object Creation, and UI Rendering. In our study, however, DrawAllocation is considered a concrete instance of a broader code pattern category we term UI Operations. More generally, each of our identified code pattern categories encompasses a broader set of real-world cases, which may be composed of different combinations of contributing factors. Many specific patterns reported in prior research can be viewed as subtypes under these broader categories. We present six representative categories of common code patterns below.
@@
API Misuse. API misuse presents a significant concern in relation to performance issues, encompassing misuses such as incorrect API invocation, misuse of API orders, and misuse of API parameters. These errors can result in severe performance consequences, such as ANR and delays. For example, passing a null context (a bad practice in parameter value passing) to PendingIntent may result in an Application Not Responding (ANR) error. This pattern may involve multiple contributing factors. Specifically, if the context is derived from an Activity or Service with an invalid lifecycle state, it can prevent the intent from being properly dispatched or executed, thereby blocking the Main Thread and eventually triggering an ANR. Besides, another common pattern involves the misuse of API order, as demonstrated in the code snippet 1 from an app com.mattermost.rnbeta, where dismissBottomSheet (line 4) is invoked before Clipboard.setString(postMessage) (line 5). In this scenario, Clipboard.setString has to wait until dismissBottomSheet is executed, leading to a delay in response time. This case mainly involves the Main Thread and UI Rendering factors.